ABC, Inc. is expected to pay an annual dividend per share of $3.50, and investors require a rate of return on S&P500 index is 10%, with the T-bill rate at 5%. What should be the current share price of ABC's stock if ABC's beta is 1.6, with a growth rate of 6%? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Share Price = ?
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

=5+1.6*(10-5)

=13%

Current price=D1/(Required return-Growth rate)

=3.5/(0.13-0.06)

=$50